OpenStudy (anonymous):

show steps in coverting 1.3 in a fraction

OpenStudy (anonymous):

well I guess you could say that .3 means 3 tenths which is 3/10. then you could look and say that the 1 would add ten more, which would make it 10+3/10= 13/10

OpenStudy (anonymous):

or simply 1 (3/10)

OpenStudy (bahrom7893):

OR even simpler: Multiply by a 10 and divide by a 10 1.3*10/10=13/10